Does the series (-1) (n + 2) ? converge absolutely, converge conditionally, or diverge? (5n) Choose the correct answer bel
Answer #1

Leibnitzs Rule: - 8 (-ig an series a is dercasing Convergent it sequence Lan> - lim q = 0 n too I otherwise 8 (-17h an diverĘ (-m (m + 2) E (-1) ain msi __ (5n, n an (2+2) ---------() a - 3 a 2 = 9 = = So, sequence and is decreasing Llim ana lim (nSo, here an= m+ 2 3 (sn) sy = na mtz him (as) ne thing motos china [ + 3n7 es el no camino Las the most beast Here e = 3 4 S